# FLARE: Faithful Logic-Aided Reasoning and Exploration

### v4

@@ 3.2 Detecting Reasoning Inconsistencies
$\forall i,\exists j\quad\text{such that}\quad\mathcal{F}_{\text{code}}^{i}=\mathcal{F}_{\text{search}}^{j}\quad\text{and}\quad\forall v,\exists q\quad\mathcal{R}_{\text{code}}^{v}=\mathcal{R}_{\text{search}}^{q}$ (5)
$\forall j,\exists i\quad\text{such that}\quad\mathcal{F}_{\text{code}}^{i}=\mathcal{F}_{\text{search}}^{j}\quad\text{and}\quad\forall q,\exists v\quad\mathcal{R}_{\text{code}}^{v}=\mathcal{R}_{\text{search}}^{q}$ (6)
With this framework in mind, we define two reasoning failure modes.
Table 2: Comparison of results across datasets for ChatGPT (gpt-3.5-turbo) and GPT-4 (gpt-4o) using Standard, CoT, Logic-LM, FLARE, and $\text{FLARE}_{SR=2}$ approaches. SR=2 refferes to a maximum of $2$ iterations of code code self-refinement.
In the first failure mode, given that some fact or relation was used in the simulated search but did not exist in the generated code, i.e. $\exists j\text{ such that }\mathcal{F}_{\text{search}}^{j}\notin\mathcal{F}_{\text{code}}$, we claim that the LLM has hallucinated. We postulate that the model either produced incomplete knowledge during formalisation to code or created a piece of non-existing information during the search. We do not consider facts that emerged during a direct inference step within the simulated search during our calculation. For example, if we are dealing with a mathematical query $4\cdot(5+6)=?$, the search would involve separately evaluating the expression $5+6=11$. In this case, $11$ will not be treated as a hallucinated fact within the search but rather as an emergent fact obtained from direct inference. The second failure mode is the reciprocal case, where a fact or relation present in the code is not used during the search. We refer to this phenomenon as sub-optimal reasoning as it shows that the LLM could not explore the problem space completely or injected unsuitable knowledge during formalisation into code.
@@ 3.3 Measuring Faithfulness
We propose a method to measure the faithfulness of the LLM reasoning process when using FLARE. As mentioned in Section 3.1, for each query in a dataset $\mathcal{D}=[\mathcal{Q}_{1},\dots,\mathcal{Q}_{|\mathcal{D}|}]$, we generate a set of codes $\Phi=[\mathcal{C}_{1},\dots,\mathcal{C}_{|\Phi|}]$ and simulated problem space searches $\Psi=[\mathcal{S}_{1},\dots,\mathcal{S}_{|\Psi|}]$. We use the Prolog engine to execute all of the codes $\Phi$ and obtain a set of correctly written programs $\Phi^{\prime}$ and exact search paths $\Psi^{\prime}$. As we do not require explicit programmatic correctness during inference in FLARE for any code $\mathcal{C}_{i}$, some Prolog executions resulting in an error are filtered out in $\Psi^{\prime}$. To assess model reasoning faithfulness towards code formalisations, we compare the search paths $\Phi^{\prime}$ obtained from Prolog execution with their designated counterparts $\Phi_{\text{\emph{gen}}}^{\prime}$ generated by the LLM from the same code. We use ROUGE (Lin, 2004) to compute the matching score for each executed and simulated search path. In particular, we use ROUGE-Lsum, which uses the longest common subsequence (LCS) over each line to obtain the final score. This method fits our cause as a line in a Prolog search execution represents a single logic step within the traversal. This allows us to measure the similarity of the reasoning contents and structure in exact and simulated searches.
@@ 4 Experimental Setup
4.1 Datasets
To evaluate FLARE, we use a benchmark of $9$ tasks covering Math Word Problems (MWP), multi-hop QA and relation inference, and $3$ common logical reasoning datasets. For testing numeric and mathematical reasoning, we follow CoT (Wei et al., 2022) by including GSM8K (Cobbe et al., 2021), SVAMP (Patel et al., 2021), MultiArith (Roy & Roth, 2015), ASDiv (Miao et al., 2020) and AQuA (Ling et al., 2017). Among these, GSM8K, SVAMP, MultiArith and ASDiv cover elementary and middle school arithmetic word problems with a set of integers or decimals as the answer. AQuA is a multiple-choice numerical, symbolic reasoning dataset where each answer is a mathematical expression containing notations, values and expressions not defined in the query. We also test FLARE using three multi-hop QA datasets. We use StrategyQA (Geva et al., 2021), which is a boolean QA task that requires sub-goal decomposition and a multi-hop reasoning strategy to answer. The example “Do all parts of the aloe vera plant taste good?” used in Figure 1, is taken from StrategyQA. The multi-hop QA testing also includes Date and Sports Understanding, subsets of BIG-Bench (bench authors, 2023). The tasks involve inferring an exact date given some calculations in the relative time period and understanding if an artificially created sports statement is feasible. Furthermore, we assess FLARE on Relational Inference using CLUTRR (Sinha et al., 2019), which involves inferring the familial relation between two entities mentioned in a natural language description of the partial family graph. We evaluate FLARE on challenging logic inference datasets: ProntoQA (Saparov & He, 2023), AR-LSAT (Zhong et al., 2021), and LogicalDeductions from BigBench (et al., 2023), focusing on harder subsets proposed by Pan et al. (2023). These datasets, covering deductive, analytical, and logical reasoning, allow us to assess FLARE’s performance. Details, including descriptions and examples, are in Table 9 of Appendix A.
4.2 Benchmarks
We compare FLARE with CoT (Wei et al., 2022) as a prompting method that reasons using natural language chains and with F-CoT (Lyu et al., 2023) and Logic-LM (Pan et al., 2023) that formalise the query into a code and offload the reasoning to an external symbolic solver. We use Llama3.1 (8B) (Dubey et al., 2024), CmDR (30B) (Cohere, 2024), CmDR+ (100B) (Cohere, 2024) and GPT3.5 (Brown et al., 2020b) ($\geq 100B$ (Ye et al., 2023)). As the coding model OpenAI Codex (code-DaVinci-002) (Chen et al., 2021) used in F-CoT has been deprecated, we replace it with the new GPT3.5 as suggested by OpenAI and recalculate the results accordingly.
@@ 5.1 Few-shot prompting
[…]
Figure 3: The figure shows the percentage of executable code per model (right) and the accuracy of the executable code when answering the queries (left).
To evaluate FLARE, we use a set of models of varying sizes on diverse benchmarks, as defined in Section 4. We compare the performance of each model while using FLARE, CoT and F-CoT prompting. The results for F-CoT and CoT on all the models are computed using the codebase of the original study (Lyu et al., 2023). We additionally compare Logic-LM and FLARE using the logic reasoning benchmarks proposed in Pan et al. (2023).
@@
Our results, presented in Table 1, show that using FLARE allows the LLMs to achieve state-of-the-art results on $7$ out of $9$ datasets, with an average $28\%$ increase over CoT. We can see a clear trend that FLARE increases the performance compared to CoT and F-CoT for all the models of varying scales. We also see that LLMs that are not explicitly tuned for coding suffer massive degeneracies when using F-CoT. We postulate that they are unable to consistently produce executable programs that satisfy a predefined scheme in F-CoT, thus resulting in an error during execution. This further highlights the value of simulating program execution using an LLM instead of using external solvers. The results show that using FLARE yields more benefit on datasets that require longer chains of multi-hop and symbolic reasoning, like AQuA and StrategyQA. Our findings in Table 2 show that FLAREachieves state-of-the-art results on $2$ out of $3$ logic inference benchmarks with $10\%$ increase over CoT and ${}7\%$ increase over Logic-LM. Following the practice in (Pan et al., 2023, Logic-LM), we also add $2$ iterations of code self-refinement to FLARE and show that the model model is able to achieve SOTA results on all $3$ benchmarks.
@@
To understand the effect of FLARE on models tuned for coding, we use GPT3.5 (Brown et al., 2020a) as it was the OpenAI suggested succession model for Codex (Chen et al., 2021) which is used in F-CoT and possesses strong coding capabilities (Ye et al., 2023). The results in Table 1 show that using FLARE is beneficial for models that are tuned for coding and boost the accuracy with a $16\%$ increase over F-CoT and $9\%$ over CoT. The reason is that many natural language queries with non-trivial formalisations are more suited to be tackled with more commonsense soft reasoning than direct code execution. This is evident in Table 1 where FLARE and CoT are consistently better than F-CoT in StrategyQA, Sports and CLUTRR. The opposite case of numeric and algorithmic heavy reasoning tasks is also covered by FLARE as it maintains strong performance similar to F-CoT on MWP problems Table 1. Consequently, FLARE allows combining algorithmic formalisation with simulated soft-reasoning, circumventing the pitfalls of using a deterministic external solver while still producing a query formalisation and problem space traversal.
@@ 5.2 Is simulating search useful?
Table 4: The table depicts the difference in the average explored paths, hops, and fails during the reasoning process, which leads to incorrect or correct answers. The purple colour illustrates that incorrect reasoning paths have fewer explorations that led to Failed search paths.
To understand if simulating a search over the problem space is useful, we compare the performance of FLARE where we only generate the plan without the subsequent code or search components. We refer to this framework setup as plan-only, which can be seen in Figure 1 if we were to use only the plan for answer generation. We completed this ablation using CmDR, CmDR+, and GPT-3.5, and we used GSM8K, AQuA, and StrategyQA as our baselines. The results in Table 3 confirm that all of the models suffer massive performance degradation from $61.1\rightarrow 49.9$ when omitting the code and the search components of FLARE. We hypothesise that this is caused by insufficient problem space exploration when using the plan-only setting. Furthermore, we have already seen in Table 1 that in methods, like F-CoT, that do not use simulated problem space exploration for soft-reasoning and only rely on plan and code, the performance also deteriorates even resulting in a complete breakdown of reasoning over the designated datasets. This can be viewed as a constrained version of FLARE with code-only execution. Consequently, our results show that simulating problem space traversal is highly beneficial as it avoids the pitfalls posed by plan-only and code-only modes by exploring the problem space more rigorously and soft-reasoning during that traversal instead of using external solvers.

### v5

@@ 3.2 Detecting Reasoning Inconsistencies
$\forall i,\exists j\quad\text{such that}\quad\mathcal{F}_{\text{code}}^{i}=\mathcal{F}_{\text{search}}^{j}\quad$ (5)
$\text{and}\quad\forall v,\exists q\quad\mathcal{R}_{\text{code}}^{v}=\mathcal{R}_{\text{search}}^{q}$
$\forall j,\exists i\quad\text{such that}\quad\mathcal{F}_{\text{code}}^{i}=\mathcal{F}_{\text{search}}^{j}\quad$ (6)
$\text{and}\quad\forall q,\exists v\quad\mathcal{R}_{\text{code}}^{v}=\mathcal{R}_{\text{search}}^{q}$
With this framework in mind, we define two reasoning failure modes. In the first failure mode, given that some fact or relation was used in the simulated search but did not exist in the generated code, i.e. $\exists j\text{ such that }\mathcal{F}_{\text{search}}^{j}\notin\mathcal{F}_{\text{code}}$, we claim that the LLM has hallucinated. We postulate that the model either produced incomplete knowledge during formalisation to code or created a piece of non-existing information during the search. We do not consider facts that emerged during a direct inference step within the simulated search during our calculation. For example, if we are dealing with a mathematical query $4\cdot(5+6)=?$, the search would involve separately evaluating the expression $5+6=11$. In this case, $11$ will not be treated as a hallucinated fact within the search but rather as an emergent fact obtained from direct inference. The second failure mode is the reciprocal case, where a fact or relation present in the code is not used during the search. We refer to this phenomenon as sub-optimal reasoning as it shows that the LLM could not explore the problem space completely or injected unsuitable knowledge during formalisation into code.
@@ 3.3 Measuring Faithfulness
We propose a method to measure the faithfulness of the LLM reasoning process when using FLARE. As mentioned in section˜3.1, for each query in a dataset $\mathcal{D}=[\mathcal{Q}_{1},\dots,\mathcal{Q}_{|\mathcal{D}|}]$, we generate a set of codes $\Phi=[\mathcal{C}_{1},\dots,\mathcal{C}_{|\Phi|}]$ and simulated problem space searches $\Psi=[\mathcal{S}_{1},\dots,\mathcal{S}_{|\Psi|}]$. We use the Prolog engine to execute all of the codes $\Phi$ and obtain a set of correctly written programs $\Phi^{\prime}$ and exact search paths $\Psi^{\prime}$. As we do not require explicit programmatic correctness during inference in FLARE for any code $\mathcal{C}_{i}$, some Prolog executions resulting in an error are filtered out in $\Psi^{\prime}$. To assess model reasoning faithfulness towards code formalisations, we compare the search paths $\Phi^{\prime}$ obtained from Prolog execution with their designated counterparts $\Phi_{\text{\emph{gen}}}^{\prime}$ generated by the LLM from the same code. We use ROUGE (Lin, 2004) to compute the matching score for each executed and simulated search path. In particular, we use ROUGE-Lsum, which uses the longest common subsequence (LCS) over each line to obtain the final score. This method fits our cause as a line in a Prolog search execution represents a single logic step within the traversal. This allows us to measure the similarity of the reasoning contents and structure in exact and simulated searches. We have also used other string-matching techniques, all of which show the same trends; thus, we report our results with ROUGE-Lsum.
Method $\text{CmDR}_{\text{plan-only}}$ $\text{CmDR}_{\text{FLARE}}$ $\text{CmDR+}_{\text{plan-only}}$ $\text{CmDR+}_{\text{FLARE}}$ $\text{GPT-3.5}_{\text{plan-only}}$ $\text{GPT-3.5}_{\text{FLARE}}$ GSM8K 24.7 52.4 40.7 71.4 36.1 68.1 AQuA 35.0 43.7 55.1 55.9 54.3 55.1 StrategyQA 65.5 67.0 75.7 70.8 62.3 65.5
@@ 4 Experimental Setup
Datasets
To evaluate FLARE, we use a benchmark of $9$ tasks spanning Math Word Problems (MWP), multi-hop QA, relation inference, and $3$ logical reasoning datasets. For numerical and mathematical reasoning, we follow CoT (Wei et al., 2022) and include GSM8K (Cobbe et al., 2021), SVAMP (Patel et al., 2021), MultiArith (Roy and Roth, 2015), ASDiv (Miao et al., 2020), and AQuA (Ling et al., 2017). GSM8K, SVAMP, MultiArith, and ASDiv focus on elementary and middle school arithmetic with integer or decimal answers. AQuA involves multiple-choice symbolic reasoning with expressions not explicitly defined in the query. We also test FLARE on three multi-hop QA tasks. StrategyQA (Geva et al., 2021) requires boolean reasoning with sub-goal decomposition (e.g., “Do all parts of the aloe vera plant taste good?” in fig.˜1). We further use the Date and Sports Understanding subsets from BIG-Bench (bench authors, 2023), which involve temporal and feasibility-based reasoning. For relation inference, we use CLUTRR (Sinha et al., 2019), which requires deducing familial relations from partial graph descriptions in natural language. We evaluate logical reasoning using ProntoQA (Saparov and He, 2023), AR-LSAT (Zhong et al., 2021), and LogicalDeductions from BIG-Bench (et al., 2023), focusing on the challenging subsets of (Pan et al., 2023). These cover deductive, analytical, and logical tasks. Dataset details and examples are in table˜7 of section˜A.1. We also study how model size affects performance and faithfulness (section˜A.1).
Benchmarks
We compare FLARE with CoT (Wei et al., 2022), which uses natural language reasoning chains, and with F-CoT (Lyu et al., 2023) and Logic-LM (Pan et al., 2023), which formalise queries into code and delegate reasoning to external solvers. Evaluated models include Llama3.1 (8B)(Dubey et al., 2024), CmDR (30B) and CmDR+ (100B)(Cohere, 2024), and GPT-3.5 (Brown et al., 2020b) ($\geq$100B (Ye et al., 2023)). As OpenAI Codex (code-DaVinci-002) (Chen et al., 2021) used in F-CoT has been deprecated, we replace it with the new GPT3.5 as suggested by OpenAI and recalculate the results.
@@ 5.1 Few-shot prompting
To evaluate FLARE, we use a set of models of varying sizes on diverse benchmarks, as defined in section˜4. We compare the performance of each model while using FLARE, CoT and F-CoT prompting. The results for F-CoT and CoT on all the models are computed using the codebase of the original study (Lyu et al., 2023). We additionally compare Logic-LM and FLARE using the logic reasoning benchmarks proposed in Pan et al. (2023).
@@
Our results, presented in table˜1, show that using FLARE allows the LLMs to achieve state-of-the-art results on $7$ out of $9$ datasets, with an average $28\%$ increase over CoT. We can see a clear trend that FLARE increases the performance compared to CoT and F-CoT for all the models of varying scales. We also see that LLMs not explicitly tuned for coding suffer massive degeneracies when using F-CoT. We postulate that they cannot consistently produce executable programs that satisfy a predefined scheme in F-CoT, thus resulting in an error during execution. This further highlights the value of simulating program execution using an LLM instead of external solvers. The results show that using FLARE yields more benefit on datasets that require longer chains of multi-hop and symbolic reasoning, like AQuA and StrategyQA. Our findings in table˜2 show that FLAREachieves state-of-the-art results on $2$ out of $3$ logic inference benchmarks with $10\%$ increase over CoT and $7\%$ increase over Logic-LM. Following the practice in (Pan et al., 2023, Logic-LM), we also add $2$ iterations of code self-refinement to FLARE and show that the model model is able to achieve SOTA results on all $3$ benchmarks.
@@
To understand the effect of FLARE on models tuned for coding, we use GPT3.5 (Brown et al., 2020a) as it was the OpenAI suggested succession model for Codex (Chen et al., 2021) which is used in F-CoT and possesses strong coding capabilities (Ye et al., 2023). The results in table˜1 show that using FLARE is beneficial for models tuned for coding and boosts accuracy with a $16\%$ increase over F-CoT and $9\%$ over CoT. The reason is that many natural language queries with non-trivial formalisations are more suited to be tackled with more commonsense soft reasoning than direct code execution. This is evident in table˜1 where FLARE and CoT are consistently better than F-CoT in StrategyQA, Sports and CLUTRR. The opposite case of numeric and algorithmic heavy reasoning tasks is also covered by FLARE as it maintains strong performance similar to F-CoT on MWP problems table˜1. Consequently, FLARE allows combining algorithmic formalisation with simulated soft-reasoning, circumventing the pitfalls of using a deterministic external solver while still producing a query formalisation and problem space traversal.
@@ 5.2 Is simulating search useful?
[…]
To understand if simulating a search over the problem space is useful, we compare the performance of FLARE where we only generate the plan without the subsequent code or search components. We refer to this framework setup as plan-only, which can be seen in fig.˜1 if we were to use only the plan for answer generation. We completed this ablation using CmDR, CmDR+, and GPT-3.5, and we used GSM8K, AQuA, and StrategyQA as our baselines. The results in table˜3 confirm that all of the models suffer massive performance degradation from $61.1\rightarrow 49.9$ when omitting the code and the search components of FLARE. We hypothesise that this is caused by insufficient problem space exploration when using the plan-only setting. Furthermore, we have already seen in table˜1 that in methods, like F-CoT, that do not use simulated problem space exploration for soft-reasoning and only rely on plan and code, the performance also deteriorates even resulting in a complete breakdown of reasoning over the designated datasets. This can be viewed as a constrained version of FLARE with code-only execution. Consequently, our results show that simulating problem space traversal is highly beneficial as it avoids the pitfalls posed by plan-only and code-only modes by exploring the problem space more rigorously and soft-reasoning during that traversal instead of using external solvers.
@@ 5.3 Faithful Reasoning Improves Performance
Model UEF (%) in Search OR (%) UR (%) in code Correct Answers $\text{Llama-3.1-8B}_{\text{FLARE}}$ 74.14 43.65 5.73 $\text{CmDR}_{\text{FLARE}}$ 59.06 35.96 4.02 $\text{CmDR+}_{\text{FLARE}}$ 64.30 34.47 4.54 $\text{GPT-3.5}_{\text{FLARE}}$ 64.46 37.55 1.90 Avg. (Correct) 65.49 37.91 4.05 Incorrect Answers $\text{Llama-3.1-8B}_{\text{FLARE}}$ 54.69 35.04 9.28 $\text{CmDR}_{\text{FLARE}}$ 54.50 32.76 6.23 $\text{CmDR+}_{\text{FLARE}}$ 44.12 24.98 8.22 $\text{GPT-3.5}_{\text{FLARE}}$ 36.02 24.44 6.94 Avg. (Incorrect) 47.33 29.31 7.67 $\Delta$ 18.16 8.60 -3.62
Table 5: The table shows how the percentage of unique emergent facts (UEF) in search, overlapping relations (OR) between code and search, and unused relations (UR) in code impact answer correctness.
As described in section˜3, using FLARE allows us to measure the faithfulness of the LLM reasoning process by comparing the simulated problem space traversals ${\Phi}^{\prime}_{\text{\emph{gen}}}$ with actual traces $\Phi$$^{\prime}$ produced from a symbolic Prolog solver. To do this, we initially compute the percentage of syntactically correct executable code each LLM produces. We have observed that all of the models are capable of producing correct executable Prolog code in $67\%$ of cases on average and $\geq 50\%$ of cases at the very least. The complete details can be seen in the top part of fig.˜3 in section˜A.5. This shows that the simulated searches ${\Phi}^{\prime}_{\text{\emph{gen}}}$ can be considered a representative sample that will be further used to accurately measure the faithfulness of the simulated search w.r.t. the generated code. After measuring the reasoning faithfulness for each model, we want to understand what impact it has on the performance of the LLM. In fig.˜2, we segment the models w.r.t. their ROUGE-Lsum scores. The results show that model performance is strongly positively correlated with reasoning faithfulness. However, we also observe that executing semantically precise code results in an accurate answer only in $47\%$ of cases on average. Refer to the bottom part of fig.˜3 in section˜A.5 for more details. Indeed, having a simulated search trace with a ROUGE-Lsum faithfulness score of $1$, would be equivalent to simply executing the program as proposed in F-CoT. Yet we have priorly shown that F-CoT struggles with reasoning tasks that are hard to formalise and require multi-hop commonsense and soft reasoning. These two discoveries show that optimal LLM reasoning, conditioned on a search in the problem space, should be increasingly faithful toward the facts, relations and the search strategy defined within the code while simultaneously maintaining the capability for soft-reasoning along more abstractly defined concepts. Our results show that FLARE allows LLMs to maintain a similar reasoning capacity.
5.4 What is important during the search?
We also analyze the reasons which can lead to optimal reasoning within an LLM. We calculate several statistics, like the average number of explored paths and the average and total hops and failures per path, for each model during the simulated traversal. The failure in a path is an invalidation of a solution for a sub-goal explored during the search, which is used for backtracking, as explained in section˜3. Calculating these statistics is simple as the search component of FLARE, seen in fig.˜1, is a structured simulation of a Prolog trace, where each line contains a hop of reasoning inference. We split these statistics for the reasoning paths that lead to correct or incorrect outcomes. Our results in table˜4 show that LLM performance and reasoning optimality are not directly connected to the amount of explored paths or multi-hop inferences per path. We also see that traces that lead to incorrect answers have a higher number of failures per path and in total. We hypothesise that LLMs with traces that were optimal for reasoning and led to correct answers could skip exploring degenerate solutions due to strong commonsense reasoning capabilities. Further analyses focus on identifying inconsistencies and failure modes (section˜3.2). By comparing relations in code with those in search traces, we measure emergent hallucinations and unused relations, highlighting areas of sub-optimal reasoning. We also assess the uniqueness of emergent facts per inference hop, indicating the extent of problem-space exploration (table˜5). The results in table˜5 show consistently over each model that, on average, traces that lead to correct answers had a higher percentage of unique emergent facts (UEF) and overlap in the relations (OR) used between the code and search, while the portion of underutilized relations (UR) was lower. This means that optimal reasoning with an LLM requires a great degree of problem-space exploration with fewer relation hallucinations during the search and more relation utilization from the defined code. This aligns with our prior discoveries, which show a strong correlation between simulated search faithfulness towards the formalised code and model performance.
6 Conclusion
This work introduces FLARE, a novel approach for logic-aided interpretable formalisation and reasoning with simulated search over the problem space. We show that models of varying scales obtain state-of-the-art results compared to prompting paradigms like CoT, F-CoT and Logic-LM. We further pinpoint that using FLARE allows us to perform soft reasoning with simulated search, making it flexible for diverse reasoning benchmarks. We introduce a method to measure model reasoning faithfulness w.r.t. the problem formalization ingrained within FLARE. Our results show that model performance is positively correlated with the faithfulness of the reasoning process. The systematic studies of the method show the benefits of using simulated search compared to natural language reasoning and external symbolic solvers. We further show that using FLARE allows us to interpretably and rigorously detect hallucinations and sub-optimal and inconsistent reasoning patterns.
Limitations
While FLARE offers significant improvements in faithfulness and interoperability, it depends on the quality of LLM-generated plans and code; errors or omissions in formalisation can propagate through the simulated search, potentially leading to incorrect or incomplete answers. The generation of formal code and simulated search traces depends heavily on the LLM’s prompt-following ability. The simulation of code execution may not fully explore extremely large or open-ended problem spaces, and prompt sensitivity can affect search thoroughness.
Risks and Impact Statement
FLARE advances the capabilities of large language models in logical reasoning and problem-solving, with potential positive impacts on applications requiring transparent and verifiable decision-making processes. The ability of the method to formalise reasoning steps and detect inconsistencies could improve reliability in high-stakes domains like healthcare decision support, educational assessment, and automated planning systems. However, this advancement also raises important considerations — the improved reasoning capabilities could be misused to automate deceptive or manipulative argumentation, and the increased persuasiveness through logical formalisation could mask underlying biases or false premises. Additionally, while FLARE improves transparency in reasoning, it may create a false sense of rigour in cases where the underlying logic is flawed but presented in a formally convincing manner.
Acknowledgments